add, simplfy by collecting like radical terms assume that all expressions under radicals represent non negative numbers √3a+9√75a^3
Answer by dabanfield(803)

add, simplfy by collecting like radical terms assume that all expressions under radicals represent non negative numbers √3a+9√75a^3
We have then:
sqrt(3*a) + 9*sqrt(5*5*3*a*a*a) =
sqrt(3*a) + 9*sqrt[(5*5)*(3*a)*(a*a)] =
sqrt(3*a) + 9*sqrt(5*5)*sqrt(3*a)*sqrt(a*a) =
sqrt(3*a) + 9*5*a*sqrt(3*a)=
sqrt(3*a) + (45*a)*sqrt(3*a) =
(1+45a)*sqrt(3a)